OpenStudy (anonymous):

Verify the identity: [(cos2 theta/sin theta)+2sin theta]=csc theta

OpenStudy (anonymous):

Use formula cos2theta = cos^2theta - sin^2 theta

OpenStudy (abmon98):

\[\cos2\theta=1-2\sin^2\theta \]\[(1-\sin^2\theta)-\sin^2\theta/2\sin \theta \] \[1-2\sin^2\theta/\sin \theta +2\sin \theta =1/\sin \theta \] \[1-2\sin^\theta+2\sin^\theta/\sin \theta \]

OpenStudy (anonymous):

That was helpful, thank you both! @Abmon98 @tgawade

OpenStudy (abmon98):

Your most Welcome :)